OpenStudy (anonymous):

Simplify: (1/2)log[3]x-2 log[3]y+4 log[3]z Solve for x: log[3]x-log[3]2=2

OpenStudy (kropot72):

\[\frac{1}{2}\log_{3} x-2\log_{3} y+4\log_{3} z=\log_{3} \sqrt{x}-\log_{3} y ^{2}+\log_{3}z ^{4}\]=\[\log_{3} \frac{\sqrt{x}}{y ^{2}}\times z ^{4}\]

OpenStudy (kropot72):

\[\log_{3} x-\log_{3} 2=2\] \[\log_{3} \frac{x}{2}=2\] \[3^{2}=9=\frac{x}{2}\] x = 18